# Completeness of closed Kleinian flat Pseudo-Riemannian Manifolds of Signature (2,2)

### v1

@@ 3.3. Reduction in ${\operatorname{SO}}(2,2)$-case
In this subsection, we apply Proposition B to the geometric setting of primary interest, namely $({\operatorname{SO}}(2,2)\ltimes\mathbb{R}^{2,2},\mathbb{R}^{2,2})$-geometry. Throughout the paper, we denote by $\mathbb{R}^{2,2}$ the affine space $\mathbb{R}^{4}$ endowed with a bilinear form of signature $(2,2)$. The isometry group preserving this bilinear form is identified with ${\operatorname{O}}(2,2)\ltimes\mathbb{R}^{4}$; the subgroup preserving orientation is ${\operatorname{SO}}(2,2)\ltimes\mathbb{R}^{4}$, and we denote by ${\operatorname{SO}}_{0}(2,2)\ltimes\mathbb{R}^{4}$ its identity component. We begin with the following classical result about isotropic planes in $\mathbb{R}^{2,2}$.
@@
Let $\Omega$ and $\Gamma$ be as in Proposition 3.1. In this proof, we consider ${\operatorname{SO}}_{0}(2,2)$, the identity component of the linear group preserving the quadratic form of signature $(2,2)$ given by $q=dxdz+dydt$. For this quadratic form, each vector of the canonical basis $(e_{1},e_{2},e_{3},e_{4})$ of $\mathbb{R}^{4}$ is isotropic. Let us consider the Cartan subgroup of ${\operatorname{SO}}_{0}(2,2)$ given in the basis $(e_{1},e_{2},e_{3},e_{4})$ by
$A=\left\{\mathrm{diag}(\lambda,\mu,\lambda^{-1},\mu^{-1})\mid\lambda,\mu>0\right\},$
and the subset
$A^{+}=\left\{\mathrm{diag}(\lambda,\mu,\lambda^{-1},\mu^{-1})\mid\lambda,\mu\geq 1\right\}.$
Considering the maximal compact subgroup $K:=S(\operatorname{O}(2)\times\operatorname{O}(2))={\operatorname{SO}}_{0}(2,2)\cap{\operatorname{SO}}(4)$ of ${\operatorname{SO}}_{0}(2,2)$, we have a Cartan decomposition ${\operatorname{SO}}_{0}(2,2)=KA^{+}K$, in the sense of Section 3.1.
The set $A^{+}$ does not contract the isotropic plane $\mathrm{Span}(e_{1},e_{2})$ (with respect to the usual euclidean norm of $\mathbb{R}^{4}$). Let $\mathcal{F}$ be a closed ${\operatorname{SO}}_{0}(2,2)$-orbit of isotropic planes. Thus, we are in the context of Proposition B, and the proof follows. ∎
@@ 5.2. Rank one kernel
  • If $\Gamma\cap P_{0}$ is trivial, then after passing to a subgroup of index $2$ of $\Gamma$, we have $p(\Gamma)\leq{\operatorname{SL}}(2,\mathbb{R})$.
@@
If $\Gamma\cap P_{0}$ is trivial, then $s_{0}\neq 0$, and hence ${\operatorname{det}}(B)=1$ for all $B\in p(\Gamma)$, so $p(\Gamma)\leq{\operatorname{SL}}(2,\mathbb{R})$. If $\Gamma\cap P_{0}$ is nontrivial, then $s_{0}=0$, and equation (11) reduces to $Bv_{0}=v_{0}$, with $v_{0}\in P_{0}$. Up to conjugacy, we may assume that $v_{0}=T_{1}$, and hence $p(\Gamma)$ is contained in the group $B_{1}$. ∎
@@ 5.2.1. $\Gamma\cap P_{0}$ trivial
From now on, and until the end of the proof of Proposition 5.6, we work under the assumptions above. We define $L:{\operatorname{GL}}_{2}^{+}(\mathbb{R})\ltimes N\to{\operatorname{GL}}_{2}^{+}(\mathbb{R})\ltimes U$ so that $L(\Gamma)$ is the linear part of $\Gamma$. Since $\Gamma\cap P_{0}$ is trivial, Lemma 5.5 implies that $p(\Gamma)\leq{\operatorname{SL}}_{2}(\mathbb{R})$, and thus $L(\Gamma)\leq{\operatorname{SL}}_{2}(\mathbb{R})\ltimes U$ (see 3.10 for the definition of $U$). However, elements of ${\operatorname{SL}}_{2}(\mathbb{R})$ commute with $U$, and therefore $L(\Gamma)$ is contained in the direct product ${\operatorname{SL}}_{2}(\mathbb{R})\times U$. Next, we define

### v2

@@ 3.3. Reduction in ${\operatorname{SO}}(2,2)$-case
In this subsection, we apply Corollary 3.7 to the geometric setting of primary interest, namely $({\operatorname{SO}}(2,2)\ltimes\mathbb{R}^{2,2},\mathbb{R}^{2,2})$-geometry. Throughout the paper, we denote by $\mathbb{R}^{2,2}$ the affine space $\mathbb{R}^{4}$ endowed with a bilinear form of signature $(2,2)$. The isometry group preserving this bilinear form is identified with ${\operatorname{O}}(2,2)\ltimes\mathbb{R}^{4}$; the subgroup preserving orientation is ${\operatorname{SO}}(2,2)\ltimes\mathbb{R}^{4}$, and we denote by ${\operatorname{SO}}_{0}(2,2)\ltimes\mathbb{R}^{4}$ its identity component. We begin with the following classical result about isotropic planes in $\mathbb{R}^{2,2}$.
@@
Let $\Omega$ and $\Gamma$ be as in Proposition 3.1. In this proof, we consider ${\operatorname{SO}}_{0}(2,2)$, the identity component of the linear group preserving the quadratic form of signature $(2,2)$ given by $q=dxdz+dydt$. For this quadratic form, each vector of the canonical basis $(e_{1},e_{2},e_{3},e_{4})$ of $\mathbb{R}^{4}$ is isotropic. In the notation of Section 3.2.1 and Corollary 3.7, we have
$\mu({\operatorname{SO}}_{0}(2,2))=\left\{\mathrm{diag}(\lambda,\mu,\lambda^{-1},\mu^{-1})\mid\lambda,\mu>0\right\}.$
Then $p_{0}=2$. Besides, the ${\operatorname{SO}}_{0}(2,2)$-orbit of the totally isotropic $2$-plane $\operatorname{Span}(e_{1},e_{2})$ is transverse, by Proposition 3.9. Hence we are in the context of Corollary 3.7, and the proof follows. ∎
@@ 5.2. Rank one kernel
  • If $\Gamma\cap P_{0}$ is trivial, then after passing to a subgroup of index $2$ of $\Gamma$, we have $p(\Gamma)\leq{\operatorname{SL}}_{2}(\mathbb{R})$.
@@
If $\Gamma\cap P_{0}$ is trivial, then $s_{0}\neq 0$, and hence ${\operatorname{det}}(B)=1$ for all $B\in p(\Gamma)$, so $p(\Gamma)\leq{\operatorname{SL}}_{2}(\mathbb{R})$. If $\Gamma\cap P_{0}$ is nontrivial, then $s_{0}=0$, and equation (11) reduces to $Bv_{0}=v_{0}$, with $v_{0}\in P_{0}$. Up to conjugacy, we may assume that $v_{0}=T_{1}$, and hence $p(\Gamma)$ is contained in the group $B_{1}$. ∎
@@ 5.2.1. $\Gamma\cap P_{0}$ trivial
From now on, and until the end of the proof of Proposition 5.6, we work under the assumptions above. We define $L:{\operatorname{GL}}_{2}^{+}(\mathbb{R})\ltimes N\to{\operatorname{GL}}_{2}^{+}(\mathbb{R})\ltimes U$ so that $L(\Gamma)$ is the linear part of $\Gamma$. Since $\Gamma\cap P_{0}$ is trivial, Lemma 5.5 implies that $p(\Gamma)\leq{\operatorname{SL}}_{2}(\mathbb{R})$, and thus $L(\Gamma)\leq{\operatorname{SL}}_{2}(\mathbb{R})\ltimes U$ (see 3.12 for the definition of $U$). However, elements of ${\operatorname{SL}}_{2}(\mathbb{R})$ commute with $U$, and therefore $L(\Gamma)$ is contained in the direct product ${\operatorname{SL}}_{2}(\mathbb{R})\times U$. Next, we define